OpenStudy (anonymous):

NEED HELP :What is the tenth term of the geometric sequence that has a common ratio of 1/3 and 36 as its fifth term?

OpenStudy (er.mohd.amir):

r=1/3 and t(5)=36 t(n)=a * r^(n-1) find a and use formula find t(10)